OpenStudy (anonymous):

Solve: sec²θ = 2tanθ+4 over the interval [0,2pi)

OpenStudy (anonymous):

\[\sec ^{2}\theta=1+ \tan ^{2}\theta\] \[\tan ^{2}\theta - 2\tan \theta - 3=0\] solve quadratic equation for tan: you'll have: \[\tan \Theta=-1 and \tan \theta=3\] what is the value of theta it corresponds...? find it :)

OpenStudy (anonymous):

I did get that far, but don't know how to figure what it corresponds to